# Simulation-based population inference of LISA’s Galactic binaries: Bypassing the global fit

### v1

@@ II.1.2 Noise Model and Data Simulation
We consider 4 years of LISA observations at a cadence of 10 seconds. We assume stationary Gaussian instrumental noise and ignore the presence of glitches and other possible noise artifacts. The PSD of the noise $S_{n}(f)$ is described by the SciRDv1 model of LISA2018ScienceReq . However, the normalizations of the acceleration noise and the optical metrology system (OMS) noise are assumed to be known only up to factors of order unit. We use the noise PSD from the LISA mission requirements document MRDv1 as implemented in LDC2022 :
@@
We produce the frequency domain waveform of each DWD using the GPU-accelerated software package GBGPU (Katz+2022, ; katz2022gbgpu, ; Cornish_Littenberg_2017, ; GBGPU-ref4, ). Given our LISA observation scenario, the GBGPU package provides the waveform for each DWD at 128 frequencies centered about the dominant frequency. The primary bottleneck is to sum the contributions of the tens of millions of DWDs in a given population into a single frequency series of length $\sim 6\times 10^{6}$. Typical vectorized broadcasting summation fails due to the size of the data. We instead leverage atomic addition within the GPU-accelerated JIT compiler, Numba. When possible, we keep most of the data generation computation on GPU, minimizing the time-consuming data transfer between CPU and GPU. To this end, for each data channel, we use Numba to sum the tens of millions of DWD GW strains, each occupying 128 frequency bins, onto the noise spectrum realization.
Finally, as a pre-processing step for the SBI, we whiten the data by dividing out a fiducial instrumental noise PSD given by Eq 5, setting $\mathcal{A}=\mathcal{O}=1$. The total time to generate a single data realization from population to strain generation is $\sim 55$ s, as compared to over a thousand hours before our accelerated implementation. Each realization contains $\sim 6\times 10^{6}$ complex 64-bit floating-point values occupying 80MB of memory. This makes it impractical to load multiple realizations in batches into the limited volatile memory of most GPUs for training the SBI. Thus, it is necessary to compress and summarize the data.
@@ II.2 Compressed Data Summary
Data summaries, $\bm{s}(\bm{D})$, are compressed representations of high-dimensional data, $\bm{D}$, that ideally retain most of the information needed to estimate the parameters of the model $\bm{\Lambda}$ (IMNN_2018, ; SBIfrontier:Kranmer+2020, ; RobustSummaries_ModelMispec_2023, ), while greatly reducing data dimensionality, $\dim(\bm{s}(\bm{D}))\ll\dim(\bm{D})$. An ideal lossless compressed summary would lead to the same posterior estimate as the full data, $p(\bm{\Lambda}|\bm{s}(\bm{D}))=p(\bm{\Lambda}|\bm{D})$. The reduced dimensionality enables larger batches of data to be loaded into memory, resulting in faster SBI training – albeit potentially incurring some information loss.
@@
We explored conventional dimensionality reduction techniques, including Principal Component Analysis (PCA) and autoencoders, to compress high-dimensional frequency series data into informative summaries for the SBI. Although these methods are widely used and effective for capturing dominant modes of variation SBIfrontier:Kranmer+2020 , they may not always preserve domain-specific structures critical for parameter recovery. Indeed, we find that naive implementations of autoencoders and PCAs tend to focus on certain noise features (such as particularly large noise fluctuations). To address this limitation, we developed a compression algorithm tailored to the specific characteristics of the DWD population spectrum. We design our summary to retain physically meaningful features of the bump in the signal power spectrum from the tens of millions of unresolved DWDs, while also providing information on the few loud, resolvable sources.
@@
First, we simplify the complex 64-bit vector representation by concatenating the 32-bit log-absolute and real values of the data, hereon denoted as $\log_{10}|\bm{D}|$ and $\Re(\bm{D})$, respectively. We empirically find that this representation offers slightly more robust training in comparison to other variations, including the real-imaginary and absolute-phase tangent pairs, as well as just the real, imaginary, or absolute values. Interestingly, variants incorporating absolute values exhibit marked improvements in the posteriors of SBI compared to those that do not. This suggests that the underlying signal-to-noise ratio of the sources may carry more informative content for population inference than the phase information associated with individually resolvable sources. However, combining the absolute value with either the real or imaginary component yields the most effective posterior estimates, offering improved coverage and accuracy (as further elaborated in § III). This suggests that jointly leveraging both amplitude and partial phase information may capture complementary aspects of the signal relevant to inference.
@@
We apply the following compression procedure independently to each channel’s frequency series realization $\bm{D}$. We first rebin the original $6\times 10^{6}$-dimensional frequency series by averaging over non-overlapping windows of 10 consecutive frequencies, resulting in a vector of size $6\times 10^{5}$. The averaged data are further subdivided into 1024 logarithmically spaced frequency bins. We also explored larger (2048) and smaller (256) numbers of bins, as well as linear bin spacing, but found log-spaced 1024 bins a good compromise of data compression and expressivity. Within each bin, we linearly fit the values as a function of frequency, with the bin-wise slope and intercept as free parameters. We then compute the standard deviation of the residuals from this linear fit, ±$\pm\sigma_{\mathrm{i}}$, for each bin of index $i$. We identify residual outliers as those that lie beyond ±$\pm 3\sigma_{\mathrm{i}}$, and quantify their importance using their $L_{2}$ norm, calculated as the square root of the sum of their squared residuals. These outliers are likely due to loud resolvable sources, and their $L_{2}$-norm thus serves as a metric that roughly quantifies their contribution to the data. Hence, for each frequency bin the data summary has four features: the slope, the intercept, the standard deviation of the residuals, and the $L_{2}$-norm of the residual outliers.
@@ II.3 Deep Neural Network
The second component is a conditional normalizing flow to model the posterior probability density. We implement a Masked Autoregressive Flow (MAF) (MAF, ) via the nflows library (nflows, ). The data embeddings from the context encoder are used to condition the flow’s transformation of a ten-dimensional standard normal distribution into the ten-dimensional posterior density $p_{\mathrm{SBI}}(\bm{\Lambda}|\bm{s}(\bm{D}))$. We design a MAF with 16 transformations, each with 4 residual blocks with 512 hidden units. Within each residual block, densely connected layers parameterize the shift and scale functions required by the autoregressive transformation.
The deep network is trained by a cross-entropy loss that maximizes the average predicted posterior for the training dataset. We illustrate the training procedure by the yellow arrows in Fig. 1, and we summarize the architecture of the deep neural network in Tables 2, 3 of § C. We train the network on a dataset comprising $3\times 10^{6}$ pairs of $\{\bm{\Lambda},\bm{D}(\bm{\Lambda})\}$. $\bm{\Lambda}$ is sampled from a prior distribution $\Pi(\bm{\Lambda})$ that is uniform over all parameters except for the noise, which follows a Gaussian distribution as specified in §II.1.1 and §II.1.2. For each sampled $\bm{\Lambda}$, we generate ten independent realizations of the DWD population, and for each population realization, we generate ten LISA instrumental noise realizations. This approach enhances training efficiency, particularly by improving the model’s ability to distinguish between noise and signal for a given parameter. We employ a training strategy that reduces the learning rate by half if the validation loss does not decrease for more than three epochs and stops training if the loss does not improve by at least 10% over ten epochs. It takes 8 hours on an NVIDIA A100 GPU to train the network for a given channel. Inference of the posterior distribution for a given realization is produced in a few seconds.
@@ II.4 Coverage & Calibration
A probability–probability (PP) plot shows the empirical coverage against the nominal credibility levels. Ideally, a calibrated inference method yields a PP plot that lies along the diagonal line (the line of perfect calibration), indicating perfect alignment between expected and actual coverage. Deviations from this diagonal indicate under-confidence (curve above the diagonal) or over-confidence (curve below the diagonal) in the posterior distributions. We illustrate this diagrammatically in the PP plot of Fig. 1. Note that in a Bayesian PP plot (i.e., one that is averaged over the prior distribution for the parameters), the diagonal line is a necessary, but not sufficient, condition for exact coverage: some regions might overcover, while others under-cover, achieving exact coverage on average but not necessarily anywhere in parameter space. Since amortized SBI enables fast posterior computation, it becomes feasible to efficiently evaluate PP distributions and thereby rapidly assess the quality of the inferred posteriors. This would be computationally expensive and time-consuming to compute for traditional likelihood-based Bayesian samplers.
@@
We evaluate the credible regions of the ten-dimensional posteriors by computing the inverse cumulative distribution function ($F^{-1}$) from posterior samples, sorted in descending order of probability density. For a given credibility level $c_{0}$, the corresponding nominal coverage is defined as the region of parameter space containing all samples with cumulative posterior mass $c\leq c_{0}$ (the HPD region). This corresponds to a region of $\bm{\Lambda}$ satisfying $p(\bm{\Lambda}|\bm{D})\geq$ $p(F^{-1}$($c_{0}$)). An HPD region has the property of being the shortest interval (in one dimension) containing a fraction $c_{0}$ of posterior probability.
If the PP curve does not follow the ideal $45^{\circ}$ line, one can calibrate the samples from the SBI in post-processing. We calibrate the posteriors with an optimal temperature $T^{*}$ by the following expression,
@@
where $p^{*}_{\mathrm{\,SBI}}(\bm{\Lambda}|\bm{D})$ and $p_{\mathrm{\,SBI}}(\bm{\Lambda}|\bm{D})$ are the calibrated and uncalibrated SBI posteriors respectively. $T^{*}$ is the optimal temperature which minimizes the mean square difference between the ideal PP plot ($45^{\circ}$ line) and that computed by scaling the posteriors. If the uncalibrated coverage is under-confident, the SBI posteriors require a calibration temperature $T^{*}<1$ to contract the distributions. Conversely, if the SBI is over-confident, the posteriors require $T^{*}>1$ to expand the distributions. A perfectly calibrated SBI corresponds to $T^{*}=1$.
In Fig. 1, we illustrate the calculation and calibration of the empirical coverage for an example nominal coverage of 0.8. Before calibration, we illustrate an empirical coverage of 0.6, where in 3 out of 5 evaluations of the posterior probability, the ground truth $\bm{\Lambda}_{\mathrm{GT}}$ lies above the 0.8 nominal coverage contour. Upon calibrating the overconfident SBI posteriors by a temperature factor, the empirical coverage converges to 0.8 as expected. The blue and orange curves of Fig. 5 show the uncalibrated and calibrated coverages for our SBI trained on the two channels A (left) and E (right). The uncalibrated coverage for the two SBIs fall above the ideal coverage, indicating under-confidence and hence requiring an optimal temperature $T^{*}<1$ (0.62 and 0.72, respectively). We provide an example comparing the calibrated and uncalibrated posteriors for the A channel SBI in §$\S$ D. Note that after this procedure, the posteriors are calibrated on average across the prior distribution – i.e., there is no guarantee that each individual posterior will exhibit exact coverage (see e.g. Karchev:2022xyn for an example of so-called frequentist calibration, which corrects coverage at each parameter value).
@@ III Results
In Fig. 6, we show the calibrated SBI posteriors for the observation-driven DWD population model of Korol+2022 . The inset illustrates the primary mass distribution, reconstructed using the posterior samples for $m_{\mathrm{0}},\,\log_{10}m_{\mathrm{\gamma}}$ in Eq. (9). The range of values shown in Fig. 6 and all subsequent posterior figures represents the prior box. The best inferred information consists of the primary mass distribution $\mathrm{d}p\mathrm{/d}m_{1}$, the separation distribution index $\alpha$, and the total number of Galactic DWDs $\log_{10}N$. In contrast, the posterior contours of the Milky Way’s disk scale lengths are less constraining. The mild bias in the recovery of $\mathcal{A}$ is not systematic throughout the SBI, and other realizations and examples result in slightly different posteriors.
Information on the Galactic scale dimensions comes from the sky localization of the sources, whose signature resides in the phase. Since the incoherent, unresolvable sources contribute limited insight into the phase and hence spatial distribution of DWDs, the SBI instead targets the signature of resolvable sources, where the phase information is more accessible. For example, constraints on the Galactic scale height $z_{\mathrm{d}}$ have been computed for the resolvable Korol_Rossi_Barausse_2018 and unresolvable sources (Breivik+2020, ). Indeed, when passing just the log-absolute data to the summary (where the phase information is lost), the SBI failed to produce meaningful posteriors for the Galactic scale dimensions, approximately reproducing the prior. This was a primary motivation to also include the real component of the data in the summary. We further characterize information from resolvable and unresolvable sources in § IV. Fig. 6 shows that our posteriors on $z_{\mathrm{d}}$ agree with the analysis of resolvable sources in (Korol_Rossi_Barausse_2018, ). In comparison to the Galactic disk scale radius, the bulge scale radius often has a less constrained posterior, likely due to the distant location of these sources in comparison to those in the disk near the solar system. On a similar note, the bulge fraction is also not as well constrained.
@@
Fig. 6 also shows a slight positive correlation between the posteriors of $m_{\mathrm{0}}$ and $\alpha$, a negative correlation with $\log_{10}N$ (more prominent at SNRs). DWD populations from different mass distributions occupy different frequency regions (c.f. the frequency histogram in Fig. 2). Populations with a higher median mass (i.e., larger $m_{\mathrm{0}}$) have higher GW frequencies, which can result in higher SNR. Thus, increasing $\log_{10}N$, decreasing $\alpha$, and increasing $m_{\mathrm{0}}$ all increase the SNR of the sources.
SBI’s ability to perform hierarchical inference on LISA data without the need for explicit likelihood evaluations opens novel avenues of testing and refining complex binary stellar evolution models that predict the formation of DWDs. For example, based on inferences of the DWD primary mass and separation distribution, one may place constraints on stable and unstable mass transfer mechanisms. Such constraints can even help better predict the merger rates of stellar-mass binary black holes originating from isolated binary evolution Srinivasan+2023 ; Iorio+2023 . A global fit of the resolvable DWDs in LISA has already been shown to help constrain these models CE_DWDseperation_Korol+2022 ; Delfavero_Breivik+2025 . A population-level inference drawing information from orders of magnitude more sources (resolvable and unresolvable) can potentially provide more robust, statistically significant constraints and is an avenue worth exploring in future studies.
@@
In future iterations, we will explore the robustness of the data summary to different simulation models (RobustSummaries_ModelMispec_2023, ) and test the SBI for model mis-specification (AnauMontel:2024flo, ). Mismodeling can arise, for instance, from incorrect/inadequate parametric representation of the population parameters (eg., mass and separation distribution), the absence of simulating overlapping signals from other sources (like MBHBs, EMRIs, stBHBs), and from simplified assumptions on the noise (stationary, Gaussianity), which may not hold in real data.

### v2

@@ II.1.2 Noise Model and Data Simulation
We consider 4 years of LISA observations at a cadence of 10 seconds. We assume stationary Gaussian instrumental noise and ignore the presence of glitches and other possible noise artifacts. We use the noise power spectral density (PSD) from the LISA mission requirements document MRDv1 as implemented in LISA Data Challenge working group (2022). Generating realistic time-varying non-Gaussian noise currently requires end-to-end instrument simulators (Bayle and Hartwig, 2023), which are currently computationally expensive for hundreds of thousands of multi-year, 10-s-cadence runs. We therefore adopt a PSD-driven stationary noise. This is a pragmatic choice, and not a limitation of the method, because the pipeline can be re-trained with more realistic, non-stationary noise, along with other artifacts, as scalable simulators mature. However, the normalizations of the acceleration noise and the optical metrology system (OMS) noise are assumed to be known only up to factors of order unity.
@@
We produce the frequency domain waveform of each DWD using the GPU-accelerated software package GBGPU (Katz et al., 2022; Katz, 2022; Cornish and Littenberg, 2007b; Robson et al., 2018). Given our LISA observation scenario, the GBGPU package provides the waveform for each DWD at 128 frequencies centered about the dominant frequency. The primary bottleneck is to sum the contributions of the tens of millions of DWDs in a given population into a single frequency series of length $\sim 6\times 10^{6}$. The typical vectorized broadcasting method fails due to the size of the data. We instead leverage atomic addition within the GPU-accelerated JIT compiler, Numba. When possible, we keep most of the data generation computation on GPU, minimizing the time-consuming data transfer between CPU and GPU. To this end, for each data channel, we use Numba to sum the tens of millions of DWD GW strains, each occupying 128 frequency bins, onto the noise spectrum realization.
Finally, as a pre-processing step for the SBI, we whiten the data by dividing out a fiducial instrumental noise PSD given by Eq 5, setting $\mathcal{A}=\mathcal{O}=1$. [^2] Note that the noise parameters $\mathcal{A},\mathcal{O}$ are sampled in our approach. The choice $\mathcal{A}=\mathcal{O}=1$ is only adopted for the data whitening, which must be the same for all simulated/observed data.
The total time to generate a single data realization from population to strain generation is $\sim 55$ s, as compared to a few hours before our accelerated implementation. Each realization contains $\sim 6\times 10^{6}$ complex 64-bit floating-point values occupying 80MB of memory. This makes it impractical to load multiple realizations in batches into the limited volatile memory of most GPUs for training the SBI. Thus, it is necessary to compress and summarize the data. Alternatively, one would need to construct an optimal embedding representation that encapsulates the context of the data.
@@ II.2 Compressed Data Summary
Data summaries, $\bm{s}(\bm{D})$, are compressed representations of high-dimensional data, $\bm{D}$, that ideally retain most of the information needed to estimate the parameters of the model $\bm{\Lambda}$ (Charnock et al., 2018; Cranmer et al., 2020; Huang et al., 2023b), while greatly reducing data dimensionality, $\dim(\bm{s}(\bm{D}))\ll\dim(\bm{D})$. An ideal lossless compressed summary would lead to the same posterior estimate as the full data, $p(\bm{\Lambda}|\bm{s}(\bm{D}))=p(\bm{\Lambda}|\bm{D})$. The reduced dimensionality enables larger batches of data to be loaded into memory, resulting in faster SBI training – albeit potentially incurring some information loss.
@@
We explored conventional dimensionality reduction techniques, including Principal Component Analysis (PCA) and autoencoders, to compress high-dimensional frequency series data into informative summaries for the SBI. Although these methods are widely used and effective for capturing dominant modes of variation Cranmer et al. (2020), they can be difficult to train when compressing large data, as in the case of years-long LISA data. Indeed, we find that naive implementations of autoencoders and PCAs tend to focus on certain noise features (such as particularly large noise fluctuations). For PCA, achieving 95% cumulative explained variance requires more than $10^{4}$ components; using fewer components fails to recover the peak amplitudes and the overall power–spectrum shape of resolvable sources. For autoencoders, the parameter count grows rapidly with network width and depth, so the trainable encoder–decoder capacity is constrained by GPU memory, which limits reconstruction fidelity. To address these limitations, we developed a compression algorithm tailored to the specific characteristics of the DWD population spectrum. We design our summary to retain physically meaningful features of the bump in the signal power spectrum from the tens of millions of unresolved DWDs, while also providing statistical information on the few loud, resolvable sources.
@@
First, we simplify the complex 64-bit vector representation by concatenating the 32-bit log-absolute and real values of the data, hereon denoted as $\log_{10}|\bm{D}|$ and $\Re(\bm{D})$, respectively. We empirically find that this representation offers slightly more robust training in comparison to other variations, including the real-imaginary and absolute-phase tangent pairs, as well as just the real, imaginary, or absolute values. Interestingly, variants incorporating absolute values exhibit marked improvements in the posteriors of SBI compared to those that do not. This suggests that the underlying signal-to-noise ratio of the sources may carry more informative content for population inference than the phase information associated with individually resolvable sources. Alternatively, as our summaries are binned, the averaging may cause us to lose phase information while retaining amplitude. We leave this for future investigation. However, combining the absolute value with either the real or imaginary component yields the most effective posterior estimates, offering improved coverage and accuracy (as further elaborated in § III). This suggests that jointly leveraging both amplitude and partial phase information may capture complementary aspects of the signal relevant to inference.
@@
We apply the following compression procedure independently to each channel's frequency series realization $\bm{D}$. We first rebin the original $6\times 10^{6}$-dimensional frequency series by averaging over non-overlapping windows of 10 consecutive frequencies, resulting in a vector of size $6\times 10^{5}$. The averaged data are further subdivided into 1024 logarithmically spaced frequency bins. We also explored larger (2048) and smaller (256) numbers of bins, as well as linear bin spacing, but found log-spaced 1024 bins a good compromise of data compression and expressivity. Within each bin, we linearly fit the values as a function of frequency, with the bin-wise slope and intercept as free parameters. We then compute the standard deviation of the residuals from this linear fit, $\pm\sigma_{\mathrm{i}}$, for each bin of index $i$. We identify residual outliers as those that lie beyond $\pm 3\sigma_{\mathrm{i}}$, and quantify their importance using their $L_{2}$ norm, calculated as the square root of the sum of their squared residuals. These outliers are likely due to loud resolvable sources, and their $L_{2}$-norm thus serves as a metric that roughly quantifies their contribution to the data. Hence, for each frequency bin the data summary has four features: the slope, the intercept, the standard deviation of the residuals, and the $L_{2}$-norm of the residual outliers.
@@ II.3 Deep Neural Network
The second component is a conditional normalizing flow to model the posterior probability density. We implement a Masked Autoregressive Flow (MAF) (Papamakarios et al., 2017) via the nflows library (Durkan et al., 2020). The data embeddings from the context encoder are used to condition the flow's transformation of a ten-dimensional standard normal distribution into the ten-dimensional posterior density $p_{\mathrm{SBI}}(\bm{\Lambda}|\bm{s}(\bm{D}))$. We design a MAF with 16 transformations, each with 4 residual blocks with 512 hidden units. Within each residual block, densely connected layers parameterize the shift and scale functions required by the autoregressive transformation.
The deep network is trained by a cross-entropy loss that maximizes the average predicted posterior for the training dataset. We illustrate the training procedure by the yellow arrows in Fig. 1, and we summarize the architecture of the deep neural network in Tables 2, 3 of § C. We train the network on a dataset comprising $3\times 10^{6}$ pairs of $\{\bm{\Lambda},\bm{D}(\bm{\Lambda})\}$. $\bm{\Lambda}$ is sampled from a prior distribution $\Pi(\bm{\Lambda})$ that is uniform over all parameters except for the noise, which follows a Gaussian distribution as specified in §II.1.1 and §II.1.2. For each sampled $\bm{\Lambda}$, we generate ten independent realizations of the DWD population, and for each population realization, we generate ten LISA instrumental noise realizations. This approach enhances training efficiency, particularly by improving the model's ability to distinguish between noise and signal for a given parameter. We employ a training strategy that reduces the learning rate by half if the validation loss does not decrease for more than three epochs and stops training if the loss does not improve by at least 10% over ten epochs. It takes 8 hours on an NVIDIA A100 GPU to train the network for a given channel. Inference of the posterior distribution for a given realization is produced in a few seconds.
@@ II.4 Coverage & Calibration
A probability–probability (PP) plot shows the empirical coverage against the nominal credibility levels. Ideally, a calibrated inference method yields a PP plot that lies along the diagonal line (the line of perfect calibration), indicating perfect alignment between expected (nominal) and actual (empirical) coverage. Deviations from this diagonal indicate under-confidence (curve above the diagonal) or over-confidence (curve below the diagonal) in the posterior distributions. We illustrate this diagrammatically in the PP plot of Fig. 1. Note that in a Bayesian PP plot (i.e., one that is averaged over the prior distribution for the parameters), the diagonal line is a necessary, but not sufficient, condition for exact coverage: some regions in the parameter space might overcover (above the $45^{\circ}$ line), while others under-cover (below the $45^{\circ}$ line), achieving exact coverage on average but not necessarily anywhere in parameter space. Since amortized SBI enables fast posterior computation, it becomes feasible to efficiently evaluate PP distributions and thereby rapidly assess the quality of the inferred posteriors. This would be computationally expensive and time-consuming to compute for traditional likelihood-based Bayesian samplers.
@@
We evaluate the credible regions of the ten-dimensional posteriors by first computing the inverse cumulative distribution function ($F^{-1}$) from posterior samples, sorted in descending order of probability density. For a given credibility level $c_{0}$, the corresponding nominal coverage is defined as the region of parameter space containing all samples with cumulative posterior mass $c\leq c_{0}$. As the cumulative mass is calculated starting from the probability density maximum, this region is known as the highest-probability density (HPD) region for that credibility level $c_{0}$. Mathematically, this is a region of $\bm{\Lambda}$ satisfying $p(\bm{\Lambda}|\bm{D})\geq$ $p(F^{-1}$($c_{0}$)). An HPD region has the property of being the shortest interval (in one dimension) containing a fraction $c_{0}$ of posterior probability.
If the PP curve does not follow the ideal $45^{\circ}$ line, the SBI samples can be calibrated in postprocessing. To this end, we calibrate the posteriors with an optimal temperature $T^{*}$ by the following expression,
@@
where $p^{*}_{\mathrm{\,SBI}}(\bm{\Lambda}|\bm{D})$ and $p_{\mathrm{\,SBI}}(\bm{\Lambda}|\bm{D})$ are the calibrated and uncalibrated SBI posteriors respectively. $T^{*}$ is a scalar optimal temperature which minimizes the mean square difference between the ideal PP plot ($45^{\circ}$ line) and the PP plot computed by scaling the posteriors. If the uncalibrated coverage is under-confident (below the $45^{\circ}$ line), the SBI posteriors are too wide and require a calibration temperature $T^{*}<1$ to contract the distributions. Conversely, if the SBI is over-confident (above the $45^{\circ}$ line), the posteriors are too narrow and require $T^{*}>1$ to expand the distributions. A perfectly calibrated SBI (on the $45^{\circ}$ line) corresponds to $T^{*}=1$.
In Fig. 1, we illustrate the calculation and calibration of the empirical coverage for an example nominal coverage of 0.8. Before calibration, we illustrate an empirical coverage of 0.6, where in 3 out of 5 evaluations of the posterior probability, the ground truth $\bm{\Lambda}_{\mathrm{GT}}$ lies above the 0.8 nominal coverage contour. Upon calibrating the overconfident SBI posteriors by a temperature factor, the empirical coverage converges to 0.8 as expected. The blue and orange curves of Fig. 5 show the uncalibrated and calibrated coverages for our SBI trained on the two channels A (left) and E (right). The uncalibrated coverage for the two SBIs fall above the ideal coverage, indicating under-confidence and hence requiring an optimal temperature $T^{*}<1$ (0.62 and 0.72, respectively). We provide an example comparing the calibrated and uncalibrated posteriors for the A channel SBI in $\S$ D. Note that after this procedure, the posteriors are calibrated on average across the prior distribution – i.e., there is no guarantee that each individual posterior will exhibit exact coverage (see e.g. Karchev et al. (2023) for an example of so-called frequentist calibration, which corrects coverage at each parameter value).
@@ III Results
In Fig. 6, we show the calibrated SBI posteriors for the observation-driven DWD population model of Korol et al. (2022c). The inset illustrates the primary mass distribution, reconstructed using the posterior samples for $m_{\mathrm{0}},\,\log_{10}m_{\mathrm{\gamma}}$ in Eq. (9). The range of values shown in Fig. 6 and all subsequent posterior figures represents the prior box. The best inferred information consists of the primary mass distribution $\mathrm{d}p\mathrm{/d}m_{1}$, the separation distribution index $\alpha$, and the total number of Galactic DWDs $\log_{10}N$. In contrast, the posterior contours of the Milky Way's disk scale lengths are less constraining. The mild bias in the recovery of $\mathcal{A}$ is not systematic throughout the SBI, and other realizations and examples result in slightly different posteriors.
Information on the Galactic scale dimensions comes from the sky localization of the sources, whose signature resides in the phase. Since the incoherent, unresolvable sources contribute limited insight into the phase and hence spatial distribution of DWDs, the SBI instead targets the signature of resolvable sources, where the phase information is more accessible. For example, constraints on the Galactic scale height $z_{\mathrm{d}}$ have been computed for the resolvable Korol et al. (2019) and unresolvable sources (Breivik et al., 2020). Indeed, when passing just the log-absolute data to the summary (where the phase information is lost), the SBI failed to produce meaningful posteriors for the Galactic scale dimensions, approximately reproducing the prior. This was a primary motivation to also include the real component of the data in the summary. We further characterize information from resolvable and unresolvable sources in § IV. Fig. 6 shows that our posteriors on $z_{\mathrm{d}}$ agree with the analysis of resolvable sources in (Korol et al., 2019). In comparison to the Galactic disk scale radius $R_{d}$, the bulge scale radius $r_{b}$ often has a less constrained posterior, likely due to the distant location of these sources in comparison to those in the disk near the solar system. On a similar note, the bulge fraction is therefore not as well constrained.
@@
SBI's ability to perform hierarchical inference on LISA data without the need for explicit likelihood evaluations opens novel avenues of testing and refining complex binary stellar evolution models that predict the formation of DWDs. For example, based on inferences of the DWD primary mass and separation distribution, one may place constraints on stable and unstable mass transfer mechanisms. Such constraints can even help better predict the merger rates of stellar-mass binary black holes originating from isolated binary evolution Srinivasan et al. (2023); Iorio et al. (2023). A global fit of the resolvable DWDs in LISA has already been shown to help constrain these models Korol et al. (2022a); Delfavero et al. (2025). A population-level inference drawing information from orders of magnitude more sources (resolvable and unresolvable) can potentially provide more robust, statistically significant constraints and is an avenue worth exploring in future studies.
@@
In future iterations, we will explore the robustness of the data summary to different simulation models (Huang et al., 2023b) and test the SBI for model mis-specification (Anau Montel et al., 2025). Mismodeling can arise, for instance, from incorrect/inadequate parametric representation of the population parameters (eg., mass and separation distribution), the absence of simulating overlapping signals from other sources (like MBHBs, EMRIs, stBHBs), and from simplified assumptions on the noise (stationary, Gaussianity), which may not hold in real data.